A 2520-lb car travels up the 8-percent incline shown. The car is subjected to a 65-lb aerodynamic drag force and a 45-lb force due to all other factors such as rolling resistance. Determine the power output required at a speed of 68 mi/hr if (a) the speed is constant and (b) the speed is increasing at the rate of 0.05g.
o= 68 mishr 100 2008 Answers: (a) Constant speed: P = b) Increase in speed is 0.059: P =
Answer #1

a) o- toni ( od 0=4.57390 65+45) -> 68 milhy F> to Efx = man 2520 322 -xo F-16S+US-WSino- F = 65+15+ 2520x sin (4.5239) I= 31